# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ Abstract
This paper proposes a simple variant of ONS, called LightONS, which reduces the total runtime to $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ while preserving the optimal $O(d\log T)$ regret. For SXO, LightONS implies a stochastic optimization method with runtime $\widetilde{O}(d^{3}/\epsilon)$, thereby answering the open problem. Importantly, LightONS preserves the elegant structure of ONS by leveraging domain-conversion techniques from parameter-free online learning to introduce a hysteresis mechanism that delays expensive Mahalanobis projections until necessary. This design enables LightONS to serve as an efficient plug-in replacement of ONS in broader scenarios, even beyond regret minimization, including gradient norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning.
1 Introduction
Online Convex Optimization (OCO) provides a versatile framework for online learning, with deep connections to stochastic optimization, game theory, and information theory (Cesa-Bianchi and Lugosi, 2006; Hazan, 2016). Online eXp-concave Optimization (OXO), is an important instance, where each online loss function is exponentially concave (exp-concave), meaning that $\exp(-\alpha f(\mathbf{x}))$ is concave for some $\alpha>0$ with $f(\mathbf{x})$ denoting the loss function. Exp-concavity naturally arises in many machine learning applications such as linear/logistic regression (Foster, 1991; Vovk, 1997; Foster et al., 2018), portfolio selection (Cover, 1991), linear-quadratic regulator control (Foster and Simchowitz, 2020), and so on. From a theoretical perspective, it introduces rich structures beyond convexity, allowing algorithms to exploit the local geometry of the loss landscape, often through local norms induced along the optimization trajectory. Such structures yield sharper statistical guarantees: Exp-concave losses admit algorithms with minimax-optimal regret $O(d\log T)$ (Ordentlich and Cover, 1998), an exponential improvement over the $\Omega(\sqrt{T})$ lower bound for convex losses (Abernethy et al., 2008).
Online Newton Step (ONS) (Hazan et al., 2007) is the de facto standard for OXO, which achieves an optimal regret bound of $O(d\log T)$ with only constant time and space per round. ONS exhibits remarkable simplicity, which has driven many advances in diverse optimization settings (Cutkosky and Orabona, 2018; Orabona et al., 2012; Luo et al., 2016) and in machine learning applications even beyond regret minimization, such as generalized linear bandits (Zhang et al., 2016; Jun et al., 2017; Zhang et al., 2025). The important insight of ONS is to leverage exp-concavity by maintaining a Hessian-related matrix that captures the local geometry of the optimization trajectory, which is crucial for a balance between statistical optimality and computational practicality. Indeed, all known other optimal OXO algorithms that avoid ONS-like Hessian maintenance, such as Exponential Weight Online Optimization (EWOO) (Hazan et al., 2007), incur prohibitive $O(T^{24})$ time complexity due to integrating over the domain (Bubeck et al., 2018).
However, the time-varying Hessian-related matrix in ONS necessitates a Mahalanobis projection at each round to ensure feasibility, which introduces the computational bottleneck. Specifically, the Mahalanobis projection solves the quadratic program $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}(\mathbf{x}-\mathbf{y})^{\mathsf{T}}M(\mathbf{x}-\mathbf{y})$ for some positive-definite and symmetric matrix $M$ and convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$. When the domain $\mathcal{X}$ is bounded, this entails $\Omega(d^{\omega})$ arithmetic operations due to matrix factorizations such as matrix square root (Golub and Van Loan, 2013).  [^1] The factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. Even for simple domains such as the unit ball and the probability simplex, Mahalanobis projection requires $\widetilde{O}(d^{\omega})$ time (see Section˜3.3) and $\widetilde{O}(d^{\omega+0.5})$ time (via interior-point methods (Nesterov and Nemirovskii, 1994)), respectively. Thus, the total runtime of ONS can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. Although the matrix-multiplication exponent satisfies $\omega<2.3714$ theoretically (Alman et al., 2025), linear algebra libraries typically operate with $\omega=3$, resulting in a runtime of $\widetilde{O}(d^{3}T)$ in practice. In contrast, for convex and strongly convex online optimization, Online Gradient Descent (OGD) (Zinkevich, 2003) requires only Euclidean projections and achieves minimax-optimal regret with a runtime of $O(dT)$ for these simple domains.
[…]
Related works.
@@ 1 Introduction
The quest for computationally efficient OXO algorithms bifurcates into two main research lines. The first, which we focus on in this paper, aims to minimize runtime while preserving optimal regret. The most relevant work is by Mhammedi and Gatmiry (2023), who proposed the OQNS (Online Quasi-Newton Steps) algorithm that attains the optimal regret $O(d\log T)$ with runtime $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$. They employ a log-barrier to eliminate Mahalanobis projections, transferring the computational burden to Hessian-inverse evaluations under log-barrier regularization. However, their method breaks from canonical algorithmic frameworks such as online mirror descent (OMD) and Follow-the-Regularized-Leader (FTRL) (Hazan, 2016), making it difficult to accommodate various local norms and thereby limiting applicability beyond OXO. A detailed comparison of ONS and our method with OQNS (Mhammedi and Gatmiry, 2023) is deferred to Section˜5.
The second research line aims to reduce regret within a time or space budget. Prominent in this line are works using matrix sketching, such as (Luo et al., 2016). Their method achieves both runtime and working memory linear in $d$, albeit under additional assumptions on the loss functions and domains. Other works in this line include projection-free methods that trade statistical optimality for computational gains, leading to suboptimal regret bounds like $O(T^{2/3})$ (Garber and Kretzu, 2023; Wan et al., 2022). The trade-off between statistical performance and computational cost is a foundational theme in machine learning. Our work advances this direction from an online optimization perspective, achieving a balance among these competing goals while preserving the flexibility to exploit problems’ local geometry. Our focus on asymptotic online behaviors contrasts with progress on offline optimization problems (Servedio, 1999; Raz, 2016; Lyu et al., 2023; Sharan et al., 2019; Peng and Rubinstein, 2023).
@@
Table 1: Algorithmic upper bounds on regret and total runtime of OXO algorithms with respect to $d$ and $T$ over a simple domain of the unit ball. “ONS-like” indicates that the method can be integrated into other settings where ONS is a key tool, including gradient-norm adaptivity, parametric bandits, and memory-efficient OXO.
@@
OGD (Zinkevich, 2003) $\sqrt{T}$ dTdT
ONS (Hazan et al., 2007) $d\log{T}$ $d^{\omega}T\log T$
[…]
@@
We propose Light Online Newton Step (LightONS), a simple ONS variant that substantially reduces the total runtime while achieving optimal regret. Our method preserves the elegant structure of ONS and thus inherits its applicability across various scenarios. Our contributions are three-fold:
  • An optimal and efficient algorithm for OXO. As summarized in Table˜1, LightONS attains the best-known total runtime $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ to achieve the minimax-optimal regret $O(d\log T)$ for OXO. Crucially, LightONS matches ONS’s dependence on all problem parameters in OXO ($T$, $d$, $D$, $G$, $\alpha$), whereas the prior method (Mhammedi and Gatmiry, 2023) suffers from large multiplicative constants. Empirical validations in Section˜3.4 corroborate the theoretical superiority of our method.
  • An optimal and efficient algorithm for SXO. With the online-to-batch conversion, LightONS yields, up to poly-logarithmic factors, the optimal sample complexity $T=\widetilde{O}(d/\epsilon)$ to achieve excess risk $\epsilon$, thus reducing the total runtime to $\widetilde{O}(d^{3}/\epsilon)$. Our result answers a COLT’13 open problem posed by Koren (2013). In Section˜4.3, we provide evidence that the runtime $\widetilde{O}(d^{3}/\epsilon)$ is unlikely to be improved in practice.
  • Applicability across various scenarios. LightONS preserves the online mirror descent (OMD) framework of ONS and inherits ONS’s structural flexibility, especially in accommodating various local norms. In Section˜5, we demonstrate plugging LightONS into scenarios which extend beyond regret minimization settings, including gradient-norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning. In contrast, the prior method (Mhammedi and Gatmiry, 2023), tailored for OXO, lacks the flexibility to fit for these scenarios.
@@
The remainder of this paper is organized as follows. Section˜2 presents preliminaries. Section˜3 presents our method and key analytical ingredients. Section˜4 discusses the implications of LightONS to SXO, which answers a COLT’13 open problem. Section˜5 demonstrates the broad applicability of LightONS inherited from ONS. Section˜6 concludes the paper and discusses future directions. All omitted details of algorithms and proofs are deferred to the appendices.
@@ 2 Preliminaries
In this section, we introduce notations used throughout this paper, formalize the problem setting of Online eXp-concave Optimization (OXO), and review important progress.
@@
Let $a]_{+}=\max\left\{0,a\right\}$ be the hinge function, let $N]=\left\{1,\dots,N\right\}$ be the index set, and let $\mathcal{B}(R)=\left\{\mathbf{x}\,|\,\left\|\mathbf{x}\right\|_{2}\leq R\right\}$ be the Euclidean ball of radius $R$ centered at the origin. Let $\nabla^{2}f(\mathbf{x})^{\mathsf{T}}$ and $\nabla^{2}f(\mathbf{x})^{-1}$ denote the transpose and inverse of the Hessian matrix, and let $\left\|\mathbf{x}\right\|_{M}=\sqrt{\mathbf{x}^{\mathsf{T}}M\mathbf{x}}$ be the Mahalanobis norm induced by a positive-definite and symmetric matrix $M$. Let $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{M}^{2}$ be the Mahalanobis projection of $\mathbf{y}$ onto a compact convex set $\mathcal{X}$, which exists and is unique (Boyd and Vandenberghe, 2004), and let $\Pi_{\mathcal{X}}[\mathbf{y}]=\Pi^{I}_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$ be the Euclidean projection. For shorthand, we write $\texttt{EP}_{\mathcal{X}}$ and $\texttt{MP}_{\mathcal{X}}$ for runtime of Euclidean and Mahalanobis projection onto $\mathcal{X}$ respectively.
@@ 2.1 Problem Setting
The regularity of the domain and the loss functions is formally stated below (Hazan, 2016).
@@
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is differentiable and $G$-Lipschitz, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\nabla f_{t}(\mathbf{x})\right\|_{2}\leq G$.
Exponentially concave (exp-concave) functions are defined below (Kivinen and Warmuth, 1999; Cesa-Bianchi and Lugosi, 2006).
Definition 1 (exp-concavity).
For a compact convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and a positive real number $\alpha$, a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave if and only if $g(\mathbf{x})=\exp\left(-\alpha f(\mathbf{x})\right)$ is concave.
Online eXp-concave Optimization (OXO) follows OCO with the exp-concavity of loss functions (Hazan et al., 2007; Mhammedi and Gatmiry, 2023), as stated below.
@@
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave.
@@ 2.2 Important Progress
In this subsection, we review two important OXO algorithms, namely Online Newton Step (ONS) (Hazan et al., 2007) and Online Quasi-Newton Step (OQNS) (Mhammedi and Gatmiry, 2023).
2.2.1 Online Newton Step (ONS)
Online Newton Step (ONS) (Hazan et al., 2007) is summarized in Algorithm˜1. The following proposition states the theoretical guarantees of ONS.
Proposition 1 (Theorem 2 of Hazan et al. (2007)).
Under Assumptions˜1, 2 and 3, ONS (Algorithm˜1) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (1)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. The total runtime of ONS is
$O\left(\left(\texttt{MP}_{\mathcal{X}}+d^{2}\right)T\right)=\widetilde{O}(d^{\omega}T).$ (2)
ONS follows the classical framework of online mirror descent (OMD) (Orabona, 2019), equipped with a time-varying Mahalanobis norm induced by the Hessian-related matrix $A_{t}$. The computational bottleneck of ONS lies in Line 6 in Algorithm˜1, the Mahalanobis projection $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$. Once the decision $\widehat{\mathbf{x}}_{t+1}$ exits the domain $\mathcal{X}$, it must be projected back. Since each Mahalanobis projection requires $\Omega(d^{\omega})$ arithmetic operations and ONS projects in $O(T)$ rounds in the worst case, the crippling total runtime $\widetilde{O}(d^{\omega}T)$ emerges, which in practice evaluates to $\widetilde{O}(d^{3}T)$.
ONS and many other practical OXO algorithms are built upon a key property of exp-concavity that bridges its curvature and the Hessian-related matrix, as stated in the following lemma, which is essentially Lemma 3 of Hazan et al. (2007) with improved constants. The proof of this lemma is provided in Section˜A.2.
@@ 2.2.1 Online Newton Step (ONS)
If a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave and differentiable, then for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, $D\geq\left\|\mathbf{x}-\mathbf{u}\right\|_{2}$, $G\geq\left\|\nabla f(\mathbf{x})\right\|_{2}$, $\gamma\leq\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, it holds that
$f(\mathbf{x})-f(\mathbf{u})\leq\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2},$ (3)
or equivalently,
$f(\mathbf{u})\geq f(\mathbf{x})+\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{u}-\mathbf{x})+\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|_{\gamma\nabla f(\mathbf{x})\nabla f(\mathbf{x})^{\mathsf{T}}}^{2}.$ (4)
We remark that, with this lemma, Assumption˜1 turns crucial for achieving the optimal regret $O(d\log T)$. Because this lemma always requires an explicit diameter $D$ to determine the curvature parameter $\gamma$, even if the problem is unconstrained. Otherwise, as $D$ increases to infinity, $\gamma$ shrinks to zero and the curvature of exp-concavity vanishes.
2.2.2 Online Quasi-Newton Step (OQNS)
@@
Online Quasi-Newton Step (OQNS) (Mhammedi and Gatmiry, 2023) effectively reduces the total runtime to reach the asymptotically optimal regret $O(d\log T)$ for OXO, as stated in the following proposition.
Proposition 2 (Theorem 9 of Mhammedi and Gatmiry (2023)).
@@
Under Assumptions˜1, 2 and 3, OQNS (Algorithm 3 of Mhammedi and Gatmiry (2023)) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{5d}{\overset{\scriptscriptstyle\circ}{\gamma}}\log\left(d+T\right)+\frac{11DGd}{2}\log T+3DGd,$ (5)
where $\overset{\scriptscriptstyle\circ}{\gamma}=\frac{1}{2}\min\left\{\frac{1}{2DG},\alpha\right\}$. The total runtime of OQNS is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\log T\right)T+d^{\omega}\sqrt{T\log T}\right).$ (6)
@@
OQNS eliminates Mahalanobis projections with a log-barrier, such as $\log(1-\left\|\mathbf{x}\right\|_{2}^{2})$ for $\mathcal{X}=\mathcal{B}(1)$, shifting the computational burden to Hessian-inverse evaluations. Key components of OQNS is illustrated below, with $\mathcal{X}=\mathcal{B}(1)$ and $G=1$ as in (Mhammedi and Gatmiry, 2023).
@@
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\textrm{Approx}\left(\nabla^{2}\Phi_{t}(\mathbf{x}_{t})^{-1}\nabla\Phi_{t}(\mathbf{x}_{t})\right),~~\text{where}$ (7a)

### v2

@@ Abstract
Online eXp-concave Optimization (OXO) is a fundamental problem in online learning, where the goal is to minimize regret when loss functions are exponentially concave. The standard algorithm, Online Newton Step (ONS), guarantees an optimal $O(d\log T)$ regret, where $d$ is the dimension and $T$ is the time horizon. Despite its simplicity, ONS may face a computational bottleneck due to the Mahalanobis projection at each round. This step costs $\Omega(d^{\omega})$ arithmetic operations for bounded domains, even for simple domains such as the unit ball, where $\omega\in(2,3]$ is the matrix-multiplication exponent. As a result, the total runtime can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. This paper proposes a simple variant of ONS, called LightONS, which reduces the total runtime to $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ while preserving the optimal regret. Deploying LightONS with the online-to-batch conversion implies a method for stochastic exp-concave optimization with runtime $\widetilde{O}(d^{3}/\epsilon)$, thereby answering an open problem posed by Koren [2013]. The design leverages domain-conversion techniques from parameter-free online learning and defers expensive Mahalanobis projections until necessary, thereby preserving the elegant structure of ONS and enabling LightONS to act as an efficient plug-in replacement in broader scenarios, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO.
1 Introduction
Online Convex Optimization (OCO) provides a versatile framework for online learning, with deep connections to stochastic optimization, game theory, and information theory [Cesa-Bianchi and Lugosi, 2006; Hazan, 2016]. Online eXp-concave Optimization (OXO), is an important instance, where loss functions are exponentially concave (exp-concave), i.e., $\exp(-\alpha f(\cdot))$ is concave for some $\alpha>0$ with $f$ denoting the loss function. Exp-concavity naturally arises in many machine learning applications such as linear/logistic regression [Foster, 1991; Vovk, 1997; Foster et al., 2018], portfolio selection [Cover, 1991], linear-quadratic regulator control [Foster and Simchowitz, 2020], and so on. From a theoretical perspective, it introduces rich structures beyond convexity, allowing algorithms to exploit the local geometry of the loss landscape, often through local norms induced along the optimization trajectory. Such structures yield sharper statistical guarantees: Exp-concave losses admit a minimax-optimal regret of $O(d\log T)$ [Ordentlich and Cover, 1998], an exponential improvement over the $\Omega(\sqrt{T})$ lower bound for convex losses [Abernethy et al., 2008].
Online Newton Step (ONS) [Hazan et al., 2007] is the de facto standard for OXO, achieving an optimal regret bound of $O(d\log T)$ with constant time and space per round. ONS exhibits remarkable simplicity, which has driven many advances in diverse optimization settings [Orabona et al., 2012; Luo et al., 2016; Cutkosky and Orabona, 2018] and in machine learning applications even beyond regret minimization, such as generalized linear bandits [Zhang et al., 2016; Jun et al., 2017; Zhang et al., 2025]. The important insight of ONS is to leverage exp-concavity by maintaining a Hessian-related matrix that captures the local geometry of the optimization trajectory, which is crucial for a balance between statistical optimality and computational practicality. Indeed, all known optimal OXO algorithms that avoid ONS-like Hessian maintenance, such as Exponential Weight Online Optimization (EWOO) [Hazan et al., 2007], incur prohibitive time complexity $O(T^{24})$ due to integrating over the domain [Bubeck et al., 2018].
However, the time-varying Hessian-related matrix in ONS necessitates a Mahalanobis projection at each round to ensure feasibility, which introduces the computational bottleneck. Specifically, the Mahalanobis projection solves the quadratic program $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}(\mathbf{x}-\mathbf{y})^{\mathsf{T}}M(\mathbf{x}-\mathbf{y})$ for some positive-definite and symmetric matrix $M$ and compact convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$. When the domain $\mathcal{X}$ is bounded, best known implementations require $\Omega(d^{\omega})$ arithmetic operations due to matrix factorizations such as matrix square root [Golub and Van Loan, 2013], where $\omega\in(2,3]$ is the matrix-multiplication exponent.  [^1] The factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. Even for simple domains such as the unit ball and the probability simplex, Mahalanobis projection requires $\widetilde{O}(d^{\omega})$ time (see Section˜B.3) and $\widetilde{O}(d^{\omega+0.5})$ time (via interior-point methods [Nesterov and Nemirovskii, 1994]), respectively. Thus, the total runtime of ONS can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. Although Alman et al. [2025] establish $\omega<2.3714$ theoretically, linear algebra libraries typically operate with $\omega=3$, yielding $\widetilde{O}(d^{3}T)$ time in practice. In contrast, for convex and strongly convex online optimization, Online Gradient Descent (OGD) [Zinkevich, 2003] requires only Euclidean projections and achieves minimax-optimal regret with a runtime of $O(dT)$ for these simple domains.
@@ 1 Introduction
Related work.
@@
The pursuit for computationally efficient OXO algorithms can be divided into two main research lines. (i) The first, which we focus on in this paper, aims to minimize runtime while preserving optimal regret. The most closely related work is by Mhammedi and Gatmiry [2023], who proposed OQNS (Online Quasi-Newton Steps), achieving the optimal regret $O(d\log T)$ with runtime $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$. They employ a log-barrier to eliminate Mahalanobis projections, transferring the computational burden to Hessian-inverse evaluations under log-barrier regularization. However, their method departs from canonical algorithmic frameworks such as online mirror descent (OMD) and Follow-the-Regularized-Leader (FTRL) [Orabona, 2019], making it difficult to accommodate diverse local norms and thereby limiting its applicability beyond standard OXO. A concrete comparison of our method with OQNS is deferred to Section˜5. (ii) The second line aims to reduce regret within a time or space budget, for example via matrix sketching techniques [Luo et al., 2016]. These methods achieve linear-in-$d$ runtime and working memory, albeit under additional assumptions on the loss functions and domains. This line also includes projection-free methods that trade statistical optimality for computational gains, leading to suboptimal regret bounds like $O(T^{2/3})$ [Garber and Kretzu, 2023; Wan et al., 2022].
@@
Table 1: Algorithmic upper bounds on regret and total runtime of OXO algorithms with respect to $d$ and $T$ over a simple domain of the unit ball. “ONS-like” indicates whether the method can be integrated into other settings where ONS serves as the backbone, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO.
@@
OGD [Zinkevich, 2003] $\sqrt{T}$ dTdT
ONS [Hazan et al., 2007] $d\log{T}$ $d^{\omega}T\log T$
OQNS [Mhammedi and Gatmiry, 2023] $d\log{T}$ $d^{2}T\log T+d^{\omega}\sqrt{T\log T}$
LightONS (This Paper) $d\log{T}$ $d^{2}T+d^{\omega}\sqrt{T\log T}$
@@
We propose Light Online Newton Step (LightONS), a simple variant of ONS that substantially reduces the runtime while retaining the optimal regret. Our method preserves the elegant structure of ONS and inherits its applicability to various scenarios. Our contributions include:
  • An optimal and efficient algorithm for OXO. As summarized in Table˜1, LightONS attains the best-known total runtime $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ to achieve the minimax-optimal regret $O(d\log T)$ for OXO. In terms of regret, LightONS matches ONS’s dependence on all problem parameters in OXO ($T$, $d$, $D$, $G$, $\alpha$), whereas the prior method [Mhammedi and Gatmiry, 2023] suffers from large multiplicative constants. Empirical validations in Appendix˜A corroborate the theoretical superiority of our method.
  • An optimal and efficient algorithm for SXO. Equipped with the online-to-batch conversion, LightONS yields (up to poly-logarithmic factors) the optimal sample complexity $T=\widetilde{O}(d/\epsilon)$ for an excess risk of $\epsilon$, thus reducing the total runtime to $\widetilde{O}(d^{3}/\epsilon)$. Our result answers a COLT’13 open problem of Koren [2013]. We further provide evidence that the runtime $\widetilde{O}(d^{3}/\epsilon)$ is unlikely to be improved in practice.
  • Applicability across various scenarios. Importantly, LightONS preserves the online mirror descent (OMD) framework of ONS and inherits ONS’s structural flexibility, especially in accommodating diverse local norms. In Section˜5, we demonstrate LightONS smoothly replaces ONS and reduces runtime in various scenarios, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO. In contrast, the prior method [Mhammedi and Gatmiry, 2023], tailored for OXO, lacks the flexibility to fit for these scenarios.
@@
The remainder of this paper is organized as follows. Section˜2 presents preliminaries. Section˜3 presents our method and key analytical ingredients. Section˜4 discusses the implications of LightONS to SXO, which answers a COLT’13 open problem. Section˜5 demonstrates the broad applicability of LightONS inherited from ONS. Section˜6 concludes the paper and discusses future directions. Omitted technical details are deferred to the appendices.
@@ 2 Preliminaries
In this section, we introduce notations used throughout this paper, formalize the problem setting of Online eXp-concave Optimization (OXO), and review key relevant prior work.
@@
Let $a]_{+}=\max\{0,a\}$, $N]=\{1,\dots,N\}$, and $\mathcal{B}(R)=\{\mathbf{x}\,|\,\left\|\mathbf{x}\right\|_{2}\leq R\}$. For a positive-definite and symmetric matrix $M$, let $\lambda_{i}(M)$ be its $i$-th greatest eigenvalue and let $\left\|\mathbf{x}\right\|_{M}=\sqrt{\mathbf{x}^{\mathsf{T}}M\mathbf{x}}$ be the corresponding Mahalanobis norm. Let $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{M}^{2}$ be the Mahalanobis projection, which exists and is unique [Boyd and Vandenberghe, 2004], and let $\Pi_{\mathcal{X}}[\mathbf{y}]=\Pi^{I}_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$ be the Euclidean projection. For shorthand, we write $\texttt{EP}_{\mathcal{X}}$ and $\texttt{MP}_{\mathcal{X}}$ for the runtime of Euclidean and Mahalanobis projection onto $\mathcal{X}$, respectively.
@@ 2.1 Problem Setting
The definition of exp-concavity [Kivinen and Warmuth, 1999; Cesa-Bianchi and Lugosi, 2006] and standard regularity assumptions of OXO [Hazan, 2016; Hazan et al., 2007; Mhammedi and Gatmiry, 2023] are formally stated below.
@@
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is differentiable and $G$-Lipschitz on $\mathcal{X}$, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\nabla f_{t}(\mathbf{x})\right\|_{2}\leq G$.
@@
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave on $\mathcal{X}$, i.e., $\exp\left(-\alpha f_{t}(\cdot)\right)$ is concave on $\mathcal{X}$.
@@ 2.2 Important Progress
In this subsection, we review two important OXO algorithms, namely Online Newton Step (ONS) [Hazan et al., 2007] and Online Quasi-Newton Step (OQNS) [Mhammedi and Gatmiry, 2023].
Online Newton Step (ONS).
Algorithm˜1 summarizes the update of ONS and enjoys the following theoretical guarantees on its regret and runtime.
Proposition 1 (Theorem 2 of Hazan et al. [2007]).
Under Assumptions˜1, 2 and 3, ONS (Algorithm˜1) with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (1a)
Runtime $\leq O\left(\left(\texttt{MP}_{\mathcal{X}}+d^{2}\right)T\right).$ (1b)
ONS follows the classical framework of online mirror descent (OMD) [Orabona, 2019], equipped with a time-varying Mahalanobis norm induced by the Hessian-related matrix $A_{t}$. The computational bottleneck of ONS lies in Line 6 in Algorithm˜1, the Mahalanobis projection $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$. Once the decision $\widehat{\mathbf{x}}_{t+1}$ exits the domain $\mathcal{X}$, it must be projected back. Since each Mahalanobis projection is implemented with $\Omega(d^{\omega})$ arithmetic operations and ONS projects in $O(T)$ rounds in the worst case, the crippling total runtime $\widetilde{O}(d^{\omega}T)$ emerges, which evaluates to $\widetilde{O}(d^{3}T)$ in practice.
Hazan et al. [2007] introduce the following lemma to characterize the curvature induced by exp-concavity, which is crucial to ONS and other practical OXO algorithms. This lemma is essentially Lemma 3 of Hazan et al. [2007], with improved constants. Its proof is deferred to Section˜B.1.
@@
If a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave and differentiable, then for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, $D\geq\left\|\mathbf{x}-\mathbf{u}\right\|_{2}$, $G\geq\left\|\nabla f(\mathbf{x})\right\|_{2}$, $\gamma\leq\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, it holds that
$f(\mathbf{x})-f(\mathbf{u})\leq\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2}.$ (2)
We remark that the preceding lemma necessitates the bounded domain assumption (Assumption˜1) to achieve the optimal regret $O(d\log T)$. In particular, a finite diameter $D$ is always required to obtain a non-zero curvature parameter $\gamma$, which reflects the curvature induced by exp-concavity.
Online Quasi-Newton Step (OQNS).
@@
OQNS [Mhammedi and Gatmiry, 2023] effectively reduces the runtime to reach the optimal asymptotical regret $O(d\log T)$, as the following proposition states.
Proposition 2 (Theorem 9 of Mhammedi and Gatmiry [2023]).
@@
Under Assumptions˜1, 2 and 3, OQNS (Algorithm 3 of Mhammedi and Gatmiry [2023]) with $\gamma=\frac{1}{2}\min\{\frac{1}{2DG},\alpha\}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{5d}{\gamma}\log\left(d+T\right)+\frac{11DGd}{2}\log T+3DGd,$ (3a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\log T\right)T+d^{\omega}\sqrt{T\log T}\right).$ (3b)
@@
OQNS eliminates Mahalanobis projections with a log-barrier, such as $\log\frac{1}{1-\left\|\mathbf{x}\right\|_{2}^{2}}$ for $\mathcal{X}=\mathcal{B}(1)$, shifting the computational burden to Hessian-inverse evaluations. Key components of OQNS is illustrated below, with $\mathcal{X}=\mathcal{B}(1)$ and $G=1$ as in [Mhammedi and Gatmiry, 2023].
@@
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\textrm{Approx}\left(\left(\nabla^{2}\Phi_{t}(\mathbf{x}_{t})\right)^{-1}\nabla\Phi_{t}(\mathbf{x}_{t})\right),\penalty 10000\ \penalty 10000\ \text{where}$ (4)